EVA Report – January 31th

Crew 274 EVA Report 31-01-2023

EVA # 3

Author: Sarah E. Guthrie, Commander

Purpose of EVA: To measure speed and distance of a 30 minute walking EVA as part of a logistics research study (PI: Pender)

Start time: 10:00am

End time: 11:30am

Narrative: EVA 3 departed the habitat at 10:00am for Kissing Camel Ridge as part of an research project for logistic caches. The crew traversed the surface in the direction of Kissing Camel Ridge for 30 mins to measure distance and speed. Once crew reached the base they took a 10 minutes break before returning to the habitat. The crew returned to the habitat at 11:20am. They then walked the MDRS campus and took a visual inspection before ingressing at 11:25am.

Destination: Kissing Camel Ridge

Coordinates (use UTM WGS 84): N4249700, E518700

Participants: EVA Lead – Nick Pender (HSO), Lex Lojek (Engineer), Tyler Hines (GHO), and Tony DiBernardo (Journalist)

Road(s) and routes per MDRS Map: Kissing Camel Ridge (see GPS map provided)

Mode of travel: Walking only

Journalist Report – January 31th

Crew 274 Journalist Report 1-31-2023

Sol: 2

Author: Tony DiBernardo, Crew Journalist

Title: Dandelion Day

Crew 274 woke up this morning at 7:30am to the song “Hall of Fame” by The Script and a recorded message from Dr. Albin, emphasizing that our efforts represent a small but important step in bringing humanity to the red planet, Mars. The team made pancakes, listened to their new collaborative playlist, and started on their Emotional Recognition Test and Psychological Surveys.

EVA Activity

Today, we were able to begin Nick “X-Man” Pender’s research while out on on EVA. His experiment includes incrementally longer EVAs over the course of the mission, eventually placing caches a 2-hour hike from the hab to test their ability to withstand in a simulated Martian environment and to support the crew on a 3-hour EVA. Today’s EVAs involved 2 teams of analog astronauts, led by Pender and Bill “Titan” O’Hara, hiking 30 minutes from the hab and back. Tony “Iron Man” DiBernardo joined both EVAs, collecting footage of his crewmates traversing the simulated martian landscape. The goal for today’s EVAs was to establish a 30-minute baseline traveling speed and distance while in an EVA spacesuit. We discovered that the pace of travel is 2.3 mph with an attainable distance of 1 mile in 30 minutes. Future EVAs for the supply cache experiment will utilize this information to place a cache 2 hiking hours from the habitat. The two EVAs started and ended on time with breathtaking views, fascinating rock formations, and a continuous process to settle into our EVa suits.


After the crew woke up, Tyler “Houston” Hines recorded overnight temperature data and observational documentation which showed promise for the remainder of the study. Overall, the sets of microgreens seem to have maintained their moisture and temperature throughout the second sol. Then, Tyler took a progress report of the space-exposed LDEF tomato seeds and assisted Commander Sarah “Ceres” Guthrie in planting dandelion seeds in lunar and martian regolith. The dandelion study is for a 2nd grader at Millersville Elementary School in Maryland. She is an aspiring astronaut and was curious if dandelions could be grown on Mars. ARG-1M graciously offered to facilitate this study, and both Lunar and Martian regolith were donated to the study.


Crew Astronomers Salina “Nova” Pena and Noah “Phoenix” Loy worked from the Musk Solar Observatory today to do a test run on the equipment before collecting data for their research. They observed a few Sun spots on the surface of the Sun and set up the MDRS Robotic Observatory through Skynet to observe High Amplitude Delta Scuti (HADS) Variable Star V0799 AUR.


As midday approached, the toilet started giving off a foul smell and Mission Support was contacted, as the crew suspected the toilet had an issue. Mission Support came to the hab and assessed the situation, informing the crew engineers Lex “Kepler” Lojek and Bill “Titan” O’Hara that Mission Support had to empty the toilet. During mission training prior to ingress, there was a miscommunication regarding when the toilet requires emptying into the septic system, so Mission Support clarified the instructions.

We ended the day getting an earlier start on our end-of-day reports than we did yesterday, hoping to have more time to relax, reset, and recover from the two EVAs. Tomorrow we look forward to a cold morning, more EVAs, and loading the images of the HADS Variable Star that were taken overnight.

GreenHab Report – January 31th


GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 68.6 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 11.62 gallons

Daily water usage for research and/or other purposes: 4.22

Water in Blue Tank: 174.6 gallons

Time(s) of watering for crops: 9 am

Changes to crops: N/A

Narrative: Monitored and watered GreenHab crops during morning operations, including additional watering for the tomato plants due to continued dryness. Cleaned the interior of pots and beds from dead leaves and debris. Identified a final initial list of potential plants to harvest, with the plan to harvest beans and additional herbs the following day.

Harvest: 9 grams of dill for crew dinner

Support/supplies needed: N/A

Sol Summary Report – January 31th



Sol: 2

Summary Title: Research EVA Phase 1 Completed

Author’s name: Sarah E. Guthrie “Ceres”, Commander

Mission Status: Nominal

Sol Activity Summary: Two crews completed two separate baseline logistics EVAs. Both crews traversed on foot from the Habitat to Kissing Camel Ridge. EVA #3 departed at 10:00am and returned at 11:30pm. EVA #4 departed at 2:00pm and returned at 3:30pm. Habitat bathroom required Mission Support assistance at 5:00pm when the waste overfilled. Mission Support serviced the toilet and provide remedial training to the crew. The solar dome shutter is unoperational currently, Mission Support has been contacted and providing guidance to troubleshoot the problem manually. The crew completed personal research and reports for the rest of the evening. GHO assisted planting dandelion seeds in Lunar and Martian regolith for a community submission study from a 2nd grader in Millersville, Maryland.

Look Ahead Plan: EVA #5 plans to place the first logistic cache which will be a part of a larger logistics study to continue the rest of the mission. The crew also plans to use the solar telescope to facilitate a study for Delta 2, USSF and 557th Space Weather Wing, USAF.

Anomalies in work: None

Weather: Sunny, highs -6C, low -11C

Crew Physical Status: Optimal

EVA: EVA 3 and EVA 4 completed logistic cache baseline study on foot at Kissing Camel Ridge successfully.

Reports to be filed: Sol Summary, Operations Report, Green Hab Report, Journalist Report, and Astronomer Report.

Support Requested: None

Operations Report – January 31th



Crew 274 Operations Report 1/31/2023

SOL: 2

Name of person filing report: Alexis Lojek

Non-nominal systems: Oven light, crew toilet

Notes on non-nominal systems: Oven light still needs a new bulb. We previously reported that the specifications of the oven light bulb type were emailed to mission support by crew 272 engineer. Around 1 pm the toilet started giving off a foul smell and Mission Support was contacted, as the crew suspected the toilet had an issue. Mission Support came to the habitat and assessed the situation, informing the crew engineers Lojek and O’Hara that Mission Support had to empty the toilet. There was a miscommunication regarding when the toilet requires emptying into the septic system, so Mission Support clarified what the correct answer is. We would recommend a drawing or diagram and more clear language be added to the Mission Handbook regarding the function of the habitat’s toilet. We appreciate the prompt response and clarification of toilet procedures. The toilet is now fixed and working nominally.


Spirit rover used: No

Hours: 207.5

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: 112.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: 219.5

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: No

Hours: 254.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

General notes on rovers: None of the four rovers were used on EVA. They are all charging from last night.

Summary of Hab operations: Habitat is now in nominal condition.

WATER USE: 29 gallons

Water (static tank): 416 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: During the first EVA of the day, suits 7, 8, 9, & 10 were used. During the second EVA of the day, suits 2, 3, & 7 were used. Helmets were cleaned before, between, and after EVAs. All radios were charged after mission completion and taken off the chargers after the red charging lights turned off.

Summary of GreenHab operations: GHO Hines watered plants this morning at 0846-0910 am.

WATER USE: 14 gallons used.

Heater: On

Supplemental light: 4 hours

Harvest: None

Summary of ScienceDome operations: GHO Officer Tyler Hines completed initial setup of research equipment. Experiment is being monitored.

Dual split: On.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The Musk Solar Observatory had issues in the afternoon (about 3 pm). The “black box” would not open the shutter. The Crew Astronomers included more detail in their own report. We would appreciate any advice on troubleshooting, as we plan to take a look at it tomorrow.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: The propane alarm went through its full startup cycle when power swapped over from the generator to the solar panels in the morning around 7 am. Is this something to be expected? Is this something we should be concerned about in any way (during power swap-over)? (Dr. Rupert responded to this question and stated that it was likely due to slow power turnover, will let know if it happens again). Copy on choosing a suit and using that the whole time.

Copyright © The Mars Society. All rights reserved. | Main Site